# Revisiting Who’s Harry Potter: Towards Targeted Unlearning from a Causal Intervention Perspective

### v1

@@ 3.3 A Causal Intervention Framework for Targeted Unlearning
Figure 3: Causal graph for data generation process.
@@
It is worth noting that this generation process also involves other concepts, e.g., other people and cities. However, their knowledge will remain fixed to the factual knowledge throughout the process, we thus leave them out of our causal model.
Under this perspective, our unlearning algorithm boils down to recovering the direct path between $\bm{X}$ and $Y$ and setting it as the teacher distribution, which becomes the standard deconfounding problem and will be discussed in the following.
@@ 3.4 Deriving the Teacher Distribution
Eq. (1) requires summing over output distributions governed by all instances of $E$, including both factual and counter-factual instances. However, we only have access to an LLM trained with factual knowledge $e_{0}$. Formally, we have $p_{\bm{\theta}}(Y|\bm{X}=\bm{x})\approx p(Y|\bm{X}=\bm{x},E=e_{0})$, where $p_{\bm{\theta}}$ denotes the output distribution of our LLM. How can we estimate $p(Y|\bm{X}=\bm{x},E=e)$ with counter-factual $e$’s?
@@
Step 3: In all the output instances of $Y^{\prime}$, change any mention of the name of $c^{\prime}$ back to $c$, i.e., $Y=\texttt{NameChange}(Y^{\prime},c^{\prime}\rightarrow c)$. This is achieved by moving the probability mass on the name of $c^{\prime}$ in the output distribution to the name of $c$. Appendix C discusses more implementation details.
It is worth mentioning that step 3, which is missing in WHP, is essential for accurately recovering the counter-factual distribution $p(Y|\bm{X}=\bm{x},E=e)$, because this distribution only involves changing the knowledge of the person, not changing the person identity. In other words, when generating a passage for Wattenbach, we want the passage to talk about the same person with alternative knowledge, but not changing the subject to a different person. As discussed in Appendix F, Step 3 is essential for avoiding mistakes of sudden subject changes.
@@ 3.5 Training a Student LLM
where $\mathcal{D}$ represents the documents used for training, and $\bm{x}$ is sampled from each position in the documents. The standard version of our method uses the provided unlearning documents as $\mathcal{D}$, e.g., Wiki pages of the unlearning targets. We also explore training on fictitious documents containing non-factual information about the target, to demonstrate the possibility to unlearn without accessing user’s factual information (details in Appendix D).
@@ 3.6 Connection to Who is Harry Potter
Counter-factual prompting. In Step 2 of section 3.4, we add an explicit prompt asking the LLM to use the replacement entity’s knowledge. This is important when the input contains conflicting facts after the name change. As shown in Appendix F, this design improves unlearning performance.
@@ 4.1 Dataset Construction
WPU contains a set of persons as unlearning targets, their associated unlearning documents, and test data in a free-response question-answering (QA) format to evaluate three types of knowledge. ❶ Forget QA covers information about the unlearning targets mentioned in unlearning documents, e.g., Q: ‘What position did Wilhelm Wattenbach hold at Berlin?’ A: ‘Professor of history’ for the target Wattenbach. ❷ Hard-retain QA covers unrelated information about other entities mentioned in unlearning documents, e.g., the city of Rantzau on Wattenbach’s Wiki page. ❸ General-retain QA covers information about unrelated persons, e.g., Elon Musk. We will describe the construction of each part below, with more details in Appendix A.
@@ 4.2 Forgetting Persons
Metrics. Table 1 defines the five criteria for the targeted unlearning task, which are measured by the following metrics (details in Appendix B). ❶ ROUGE calculates the ROUGE-L score Lin (2004) between ground-truth (GT) and generated answers. Since GT answers in our dataset are concise, ROUGE evaluates the correctness of generated answers. ❷ GPT privacy score: Given the question, GT answer, and model-generated response, GPT-4 rates how well the response protects the unlearning target’s factual information, with scores from $\{1,2,3\}$, where $3$ indicates no factual leakage. ❸ GPT quality score: Given the question and generated response, GPT-4 assigns scores from $\{1,2,3\}$ to evaluate response quality, where $3$ denotes fluent, relevant, and appropriate responses, regardless of correctness. ❹ Rep-4 Welleck et al. (2020) measures the portion of duplicate 4-grams in a generated response. ❺ GPT rejection rate calculates the percentage of responses that reject the question by indicating the information is unavailable (e.g., the person does not exist or cannot be recalled). [^3] A response that does not reject the question can be either hallucination or leakage of factual information, but a high rejection rate prevents both cases. With these metrics, normalized to $[0,1]$, the five criteria are evaluated as in Table 1.
@@
Baselines. We compare seven baselines (details in Appendix C): ❶ Gradient ascent (GA) Yao et al. (2024b) maximizes cross-entropy loss on unlearning documents. ❷ Negative preference optimization (NPO) Zhang et al. (2024) modifies GA into a bounded loss to prevent model degeneration. Both GA and NPO include a regularization term minimizing cross-entropy loss on Wiki pages of 100 persons not in the test data. ❸ Prompt Lynch et al. (2024); Thaker et al. (2024) prompts the LLM to not generate anything related to the unlearning targets. ❹ Prompt-distill uses outputs of ❸ as a teacher and trains an LLM to mimic teacher responses on additionally created QA pairs about the unlearning targets. Since most teacher responses are like ‘I don’t know’, ❹ resembles works that explicitly train the LLM to generate such responses Ishibashi and Shimodaira (2024); Maini et al. (2024). To prevent the LLM from refusing all questions, we add a term training the LLM to correctly answer unrelated questions. ❺ Deliberate imagination (DI) Dong et al. (2024) uses the LLM’s output distribution on original unlearning documents as the teacher but reduces the logit of the original token by a constant. ❻ WHP in Eldan and Russinovich (2023). Since their code is not available, we re-implement it based on our understanding of the method. ❼ WHP+ (Ours-1), which is an instance of our framework where all improved designs in section 3.6 are included except for aggregating multiple distributions. In short, ❶, ❷, and ❹ require additional retain documents, and ❹ further converts them to QA pairs. We test all methods on Llama2-7b-chat Touvron et al. (2023).
@@
Implementation details. We train the model on unlearning documents (except two prompt-based methods) and evaluate it on the three QA sets in WPU. For our method, the teacher aggregates 20 distributions (replacement names in Appendix C).
@@
Results. Figure 4 shows the results on forgetting 2 and 100 persons (full results in Appendix D). We report the average of 5 different sets of 2 persons. Each criterion is normalized by the maximum across all methods, so the highest score is 100.
@@
There are five observations. First, our method achieves high performance in all criteria, whereas baselines fall short in some. For example, GA has low response quality, often generating gibberish. Its model utility also degrades, as it trains on the entire document without differentiating information to retain or forget. The two prompt-based methods achieve high unlearning efficacy but perform poorly under adversarial attacks, indicating the knowledge is not truly removed. The LLM also incorrectly refuses unrelated questions, which hurts model utility. Second, without accessing any retain documents, our method sustains a high model utility, verifying that our causal intervention framework only perturbs the unlearning target’s knowledge. Third, while we do not explicitly optimize for fewer hallucinations, our method responds to over $90\%$ questions by indicating that the information is unavailable. In section 4.4, we show that aggregating multiple distributions is critical for this behavior. Fourth, Ours-1 significantly outperforms WHP, demonstrating the benefits of better designs informed by our framework. Finally, comparing Ours-1 and Ours, we observe that aggregating multiple distributions effectively reduces the hallucination rate, especially in the forget 2 persons setting. A more in-depth study is presented in section 4.4. Table A6 shows sample outputs verifying these observations.
@@
Figure 5: Forget Quality ($\uparrow$) vs. Model Utility ($\uparrow$) on TOFU (average of $3$ seeds). For clarity, values above $0.1$ are in linear scale and those below $0.1$ are in log scale.
@@ 4.3 Forgetting Authors and Books
Adaptation for WHP. We add an important design to improve WHP on TOFU. We treat authors and books as unlearning targets and replace their names during teacher construction. The original WHP does not train the student LLM on tokens within a name span. However, a model does not know the author or the book should assign low probabilities to its name. Based on our framework, we can achieve this by constructing the teacher given perturbed prefix of the name, e.g., predicting the last name given a different first name (details in Appendix C). The model with this modification and our other designs (except for aggregating multiple distributions) is denoted as WHP+ (Ours-1).
Results. Figure 5 shows the results on forgetting $10\%$ authors (full results in Appendix E). Following Zhang et al. (2024), we evaluate models after every epoch and report the epoch with the best forget quality. An ideal method should be in the top-right corner. There are two observations. First, our two methods achieve the best forget quality and a high model utility, without access to any retain data. Most baselines, including WHP, fail to achieve a $p$-value higher than 0.05, indicating unsuccessful unlearning. Second, Ours-1 better approximates the retrained model than Ours. Analyses in Appendix E show that as more distributions are aggregated, the student LLM has a flatter output distribution, where the knowledge being unlearned and its perturbations have similar probabilities, but the retrained model has more spiky distributions. These results, together with section 4.4, show that our framework can trade off between various criteria. On the one hand, aggregating more distributions leads to desirable behaviors such as fewer hallucinations. One the other hand, using one distribution better approximates a retrained model.
@@ 4.4 Ablation Study
Figure 6 (a) shows hallucination avoidance as a function of $N$. We report the performance of directly using the teacher distribution to answer questions (in red), as well as the student model (in blue). Notably, increasing $N$ reduces hallucinations for the teacher distribution. As aggregating multiple names flattens the output distribution, responses like ‘I don’t known’ emerge. The student model, which is trained only on the Wiki pages, generalizes this behavior to the QA format. Figure 6 (b) shows the overall performance of the student model, which illustrates that increasing $N$ leads to better performance and more stable training target, as shown by the fewer outliers. The benefits of our other designs are shown in Appendix F.
@@ 6 Acknowledgements
The work of Yujian Liu and Shiyu Chang was partially supported by National Science Foundation (NSF) Grant IIS-2207052, NSF Grant IIS-2302730, and the UC Santa Barbara IEE IGSB SW Impact Grant. The computing resources used in this work were partially supported by the Accelerate Foundation Models Research program of Microsoft. Tommi Jaakkola acknowledges support from the MIT-IBM Watson AI Lab and the NSF Expeditions grant (award 1918839: Collaborative Research: Understanding the World Through Code).
@@ 7 Limitations
There are two limitations in our work that can be further improved. First, neither our method nor the evaluated baselines provide a theoretical guarantee of unlearning of the target knowledge. Instead, we measure the performance of all methods under adversarial attacks to empirically evaluate the worst-case unlearning performance. Therefore, the conclusions drawn in this paper pertain specifically to the two jailbreak attacks being considered Anil et al. (2024); Schwinn et al. (2024). We encourage future works to expand our evaluations of the unlearned model. Second, although our method maintains high model utility compared to baselines, there is still some degradation in utility compared to the original model. This degradation may result from the complex interactions between various knowledge in the LLM. Future works can explore other methods to better maintain model utility, such as surgically modify model parameters instead of full fine-tuning Lee et al. (2023).
@@ 8 Ethical Considerations and Use of Data
Appendix A Construction of WPU

### v2

@@ 3.3 A Causal Intervention Framework for Targeted Unlearning
Figure 3: Causal graph for the data generation process.
@@
It is worth mentioning two important assumptions that make our structural causal model valid:
Assumption 1: Pre-assumed causal relations. We construct the causal graph using pre-assumed causal relations between the random variables based on our prior knowledge. However, the causal relations may not hold in certain cases. For example, consider the input $\bm{X}=$Germany is the birth country of’ and $Y=$Wattenbach’. In this case, the fact that $Y=$Wattenbach’ likely decides $\bm{X}$ mentions Germany instead of other countries, indicating a causal edge from $Y$ to $\bm{X}$. Fortunately, for most unlearning documents considered in this paper, the reversing direction would not occur (e.g., most Wikipedia sentences begin with the name of the unlearning target). When it does occur, our algorithm enables a mitigation mechanism, which will be discussed in Appendix E.1.
Assumption 2: Constant remaining entities. There may be many paths connecting $\bm{X}$ and $Y$ in the causal graph, which correspond to the knowledge of other entities, e.g., other people and cities. However, given an unlearning target, we assume the knowledge of all other entities are fixed to their realizations in our current world, thus their effects can be considered as absorbed in the direct path from $\bm{X}$ to $Y$.
Under this causal perspective, our unlearning algorithm boils down to recovering the direct path between $\bm{X}$ and $Y$ and setting it as the teacher distribution, which becomes the standard deconfounding problem and will be discussed in the following.
@@ 3.4 Deriving the Teacher Distribution
Note that to apply the backdoor theorem, it is important for assumption 2 to hold, which ensures that the unlearning target’s knowledge $E$ blocks all backdoor paths from $\bm{X}$ to $Y$. Alternatively, we can cast the left-hand side of Eq.(1) as the intervention distribution conditional upon the remaining entities fixed to their real-world realizations. Appendix A elaborates this interpretation.
Eq. (1) requires summing over output distributions governed by all instances of $E$, including factual and counter-factual instances. However, we only have access to an LLM trained with factual knowledge $e_{0}$. Formally, we have $p_{\bm{\theta}}(Y|\bm{X}=\bm{x})\approx p(Y|\bm{X}=\bm{x},E=e_{0})$, where $p_{\bm{\theta}}$ denotes the output distribution of our LLM. How can we estimate $p(Y|\bm{X}=\bm{x},E=e)$ with counter-factual $e$’s?
@@
Step 3: In all the output instances of $Y^{\prime}$, change any mention of the name of $c^{\prime}$ back to $c$, i.e., $Y=\texttt{NameChange}(Y^{\prime},c^{\prime}\rightarrow c)$. This is achieved by moving the probability mass on the name of $c^{\prime}$ in the output distribution to the name of $c$. Appendix D discusses more implementation details.
It is worth mentioning that step 3, which is missing in WHP, is essential for accurately recovering the counter-factual distribution $p(Y|\bm{X}=\bm{x},E=e)$, because this distribution only involves changing the knowledge of the person, not changing the person identity. In other words, when generating a passage for Wattenbach, we want the passage to talk about the same person with alternative knowledge, but not changing the subject to a different person. As discussed in Appendix G, Step 3 is essential for avoiding mistakes of sudden subject changes.
@@ 3.5 Training a Student LLM
where $\mathcal{D}$ represents the documents used for training, and $\bm{x}$ is sampled from each position in the documents. The standard version of our method uses the provided unlearning documents as $\mathcal{D}$, e.g., Wiki pages of the unlearning targets. We also explore training on fictitious documents containing non-factual information about the target, to demonstrate the possibility to unlearn without accessing users’ factual information (details in Appendix E).
@@ 3.6 Connection to Who is Harry Potter
Counter-factual prompting. In Step 2 of section 3.4, we add an explicit prompt asking the LLM to use the replacement entity’s knowledge. This is important when the input contains conflicting facts after the name change. As shown in Appendix G, this design improves unlearning performance.
@@ 4.1 Dataset Construction
WPU contains a set of persons as unlearning targets, their associated unlearning documents, and test data in a free-response question-answering (QA) format to evaluate three types of knowledge. ❶ Forget QA covers information about the unlearning targets mentioned in unlearning documents, e.g., Q: ‘What position did Wilhelm Wattenbach hold at Berlin?’ A: ‘Professor of history’ for the target Wattenbach. ❷ Hard-retain QA covers unrelated information about other entities mentioned in unlearning documents, e.g., the city of Rantzau on Wattenbach’s Wiki page. ❸ General-retain QA covers information about unrelated persons, e.g., Elon Musk. We will describe the construction of each part below, with more details in Appendix B.
@@ 4.2 Forgetting Persons
Metrics. Table 1 defines the five criteria for the targeted unlearning task, which are measured by the following metrics (details in Appendix C). ❶ ROUGE calculates the ROUGE-L score Lin (2004) between ground-truth (GT) and generated answers. Since GT answers in our dataset are concise, ROUGE evaluates the correctness of generated answers. ❷ GPT privacy score: Given the question, GT answer, and model-generated response, GPT-4 rates how well the response protects the unlearning target’s factual information, with scores from $\{1,2,3\}$, where $3$ indicates no factual leakage. ❸ GPT quality score: Given the question and generated response, GPT-4 assigns scores from $\{1,2,3\}$ to evaluate response quality, where $3$ denotes fluent, relevant, and appropriate responses, regardless of correctness. ❹ Rep-4 Welleck et al. (2020) measures the portion of duplicate 4-grams in a generated response. ❺ GPT rejection rate calculates the percentage of responses that reject the question by indicating the information is unavailable (e.g., the person does not exist or cannot be recalled). [^3] A response that does not reject the question can be either hallucination or leakage of factual information, but a high rejection rate prevents both cases. With these metrics, normalized to $[0,1]$, the five criteria are evaluated as in Table 1. Additionally, to ensure there is no systematic bias due to the use of GPT-4 in both data generation and evaluation, we use Llama-3 Llama Team (2024) to repeat the above evaluations and observe consistent results with GPT-4’s scores (details in Appendix E.2).
@@
Baselines. We compare seven baselines (details in Appendix D): ❶ Gradient ascent (GA) Yao et al. (2024b) maximizes cross-entropy loss on unlearning documents. ❷ Negative preference optimization (NPO) Zhang et al. (2024) modifies GA into a bounded loss to prevent model degeneration. Both GA and NPO include a regularization term minimizing cross-entropy loss on Wiki pages of 100 persons not in the test data. ❸ Prompt Lynch et al. (2024); Thaker et al. (2024) prompts the LLM to not generate anything related to the unlearning targets. ❹ Prompt-distill uses outputs of ❸ as a teacher and trains an LLM to mimic teacher responses on additionally created QA pairs about the unlearning targets. Since most teacher responses are like ‘I don’t know’, ❹ resembles works that explicitly train the LLM to generate such responses Ishibashi and Shimodaira (2024); Maini et al. (2024). To prevent the LLM from refusing all questions, we add a term training the LLM to correctly answer unrelated questions. ❺ Deliberate imagination (DI) Dong et al. (2024) uses the LLM’s output distribution on original unlearning documents as the teacher but reduces the logit of the original token by a constant. ❻ WHP in Eldan and Russinovich (2023). Since their code is not available, we re-implement it based on our understanding of the method. ❼ WHP+ (Ours-1), which is an instance of our framework where all improved designs in section 3.6 are included except for aggregating multiple distributions. In short, ❶, ❷, and ❹ require additional retain documents, and ❹ further converts them to QA pairs. Additionally, we also compare with an RLHF baseline that trains the model to abstain from questions about the unlearning target, which will be discussed in Appendix E.3. The following sub-section reports the performance of all methods on Llama2-7b-chat Touvron et al. (2023). Additional results on Llama-3 Llama Team (2024) are provided in Appendix E.4.
@@
Implementation details. We train the model on unlearning documents (except two prompt-based methods) and evaluate it on the three QA sets in WPU. For our method, the teacher aggregates 20 distributions (replacement names in Appendix D).
@@
Results. Figure 4 shows the results on forgetting 2 and 100 persons (full results in Appendix E). We report the average of 5 different sets of 2 persons. Each criterion is normalized by the maximum across all methods, so the highest score is 100.
@@
There are five observations. First, our method achieves high performance in all criteria, whereas baselines fall short in some. For example, GA has low response quality, often generating gibberish. Its model utility also degrades, as it trains on the entire document without differentiating information to retain or forget. The two prompt-based methods achieve high unlearning efficacy but have low model utility, as the LLM incorrectly refuses unrelated questions. Particularly, Prompt also performs poorly under adversarial attacks, indicating the knowledge is not truly removed. Second, without accessing any retain documents, our method sustains a high model utility, verifying that our causal intervention framework only perturbs the unlearning target’s knowledge. Third, while we do not explicitly optimize for fewer hallucinations, our method responds to over $90\%$ questions by indicating that the information is unavailable. In section 4.4, we show that aggregating multiple distributions is critical for this behavior. Fourth, Ours-1 significantly outperforms WHP, demonstrating the benefits of better designs informed by our framework. Fifth, comparing Ours-1 and Ours, we observe that aggregating multiple distributions effectively reduces the hallucination rate, especially in the forget 2 persons setting. A more in-depth study is presented in section 4.4. In addition, we also evaluate the unlearned models’ generalizability to different languages and aliases of the unlearning target. Results in Appendix E.5 show that most methods are robust to such perturbations at test time. Finally, to investigate the inherent tradeoff among five criteria, we calculate the correlation between each pair of criteria and show the results in Appendix E.6. Table A6 shows sample outputs verifying the above observations.
@@
Figure 5: Forget Quality ($\uparrow$) vs. Model Utility ($\uparrow$) on TOFU (average of $3$ seeds). For clarity, values above $0.1$ are in linear scale, and those below $0.1$ are in log scale.
@@ 4.3 Forgetting Authors and Books
Adaptation for WHP. We add an important design to improve WHP on TOFU. We treat authors and books as unlearning targets and replace their names during teacher construction. The original WHP does not train the student LLM on tokens within a name span. However, a model does not know the author or the book should assign low probabilities to its name. Based on our framework, we can achieve this by constructing the teacher given perturbed prefix of the name, e.g., predicting the last name given a different first name (details in Appendix D). The model with this modification and our other designs (except for aggregating multiple distributions) is denoted as WHP+ (Ours-1).
Results. Figure 5 shows the results on forgetting $10\%$ authors (full results in Appendix F). Following Zhang et al. (2024), we evaluate models after every epoch and report the epoch with the best forget quality. An ideal method should be in the top-right corner. There are two observations. First, our two methods achieve the best forget quality and a high model utility, without access to any retain data. Most baselines, including WHP, fail to achieve a $p$-value higher than 0.05, indicating unsuccessful unlearning. Second, Ours-1 better approximates the retrained model than Ours. Analyses in Appendix F show that as more distributions are aggregated, the student LLM has a flatter output distribution, where the knowledge being unlearned and its perturbations have similar probabilities, but the retrained model has more spiky distributions. These results, together with section 4.4, show that our framework can trade off between various criteria. On the one hand, aggregating more distributions leads to desirable behaviors such as fewer hallucinations. On the other hand, using one distribution better approximates a retrained model.
@@ 4.4 Ablation Study
Figure 6 (a) shows hallucination avoidance as a function of $N$. We report the performance of directly using the teacher distribution to answer questions (in red), as well as the student model (in blue). Notably, increasing $N$ reduces hallucinations for the teacher distribution. As aggregating multiple names flattens the output distribution, responses like ‘I don’t know’ emerge. The student model, which is trained only on the Wiki pages, generalizes this behavior to the QA format. Figure 6 (b) shows the overall performance of the student model, which illustrates that increasing $N$ leads to better performance and a more stable training target, as shown by the fewer outliers. The benefits of our other designs are shown in Appendix G.
@@ 6 Acknowledgements
The work of Yujian Liu and Shiyu Chang was partially supported by National Science Foundation (NSF) Grant IIS-2338252, NSF Grant IIS-2207052, NSF Grant IIS-2302730, and the UC Santa Barbara IEE IGSB SW Impact Grant. The computing resources used in this work were partially supported by the Accelerate Foundation Models Research program of Microsoft. Tommi Jaakkola acknowledges support from the MIT-IBM Watson AI Lab and the NSF Expeditions grant (award 1918839: Collaborative Research: Understanding the World Through Code).
@@ 7 Limitations
There are two limitations in our work that can be further improved. First, neither our method nor the evaluated baselines provide a theoretical guarantee of unlearning of the target knowledge. Instead, we measure the performance of all methods under adversarial attacks to empirically evaluate the worst-case unlearning performance. Therefore, the conclusions drawn in this paper pertain specifically to the two jailbreak attacks being considered Anil et al. (2024); Schwinn et al. (2024). We encourage future works to expand our evaluations of the unlearned model. Second, although our method maintains high model utility compared to baselines, there is still some degradation in utility compared to the original model. This degradation may result from the complex interactions between various knowledge in the LLM. Future works can explore other methods to better maintain model utility, such as surgically modifying model parameters instead of full fine-tuning Lee et al. (2023).
@@ 8 Ethical Considerations and Use of Data
Appendix A A Conditional Interpretation of the Teacher Distribution
Our derivation of the teacher distribution in Eq. (1) considers remaining entities other than the unlearning target as fixed, thus absorbing their effects in the direct path from $\bm{X}$ to $Y$. Alternatively, we can also cast the teacher distribution as the intervention distribution conditional upon the remaining entities fixed to their real-world realizations.
More specifically, we define $E_{i}$ as the knowledge of the unlearning target, e.g., Wattenbach, and $E_{-i}$ as the knowledge of all other entities, e.g., other people, places, and organizations that may or may not relate to Wattenbach. Our teacher distribution estimates $p(Y|do(\bm{X}=\bm{x}),E_{-i}=e_{-i})$, where $e_{-i}$ represents the values of other entities’ knowledge in our current world. To estimate this distribution, we again apply the backdoor theorem with adjustment set $E_{i}$, which leads to
$p(Y|$ $do(\bm{X}=\bm{x}),E_{-i}=e_{-i})$ (2)
$=\sum_{e_{i}}p(Y|\bm{X}=\bm{x},E_{i}=e_{i},E_{-i}=e_{-i})$
$\cdot p(E_{i}=e_{i}|E_{-i}=e_{-i}),$
where we estimate the first term with our name change algorithm and assume the second term to be a uniform distribution over knowledge of real-world persons. In practice, we can estimate $p(Y|\bm{X}=\bm{x},E_{i}=e_{i},E_{-i}=e_{-i})$ using a pre-trained LLM, because its pre-training corpus corresponds to the knowledge of $e_{-i}$. Note that this teacher distribution precisely describes the targeted unlearning task, where knowledge of other entities are unaffected, and we only forget the unlearning target’s knowledge.
Appendix B Construction of WPU